# Are Multimodal Embeddings Truly Beneficial for Recommendation? A Deep Dive into Whole vs. Individual Modalities

### v2

@@ Abstract.
To isolate the effect of individual modalities—text or visual—we employ a modality knockout strategy by setting the corresponding embeddings to either constant values or random noise. To ensure the scale and comprehensiveness of our study, we evaluate 14 widely used state-of-the-art multimodal recommendation models. Our findings reveal that: (1) multimodal embeddings generally enhance recommendation performance—particularly when integrated through more sophisticated graph-based fusion models. Surprisingly, commonly adopted baseline models with simple fusion schemes, such as VBPR and BM3, show only limited gains. (2) The text modality alone achieves performance comparable to the full multimodal setting in most cases, whereas the image modality alone does not. These results offer foundational insights and practical guidance for the multimodal recommendation community. We will release our code and datasets to facilitate future research.
Multimodal Recommendation, Multimodal Embeddings, MMRec, Empirical Study
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Recommender systems
1. Introduction
@@
Multimodal recommendation has emerged as a powerful paradigm in modern recommender systems, enabling the integration of diverse modalities such as text, image, video, and audio. By leveraging pretrained encoders such as Sentence-BERT, Vision Transformers (ViT), and ResNet, recent state-of-the-art models—including BM3 (Zhou et al., 2023a), FREEDOM (Zhou and Shen, 2023), and SMORE (Ong and Khong, 2025)—have demonstrated impressive performance across various recommendation tasks. These models are built on the intuitive yet widely accepted assumption that incorporating multimodal embeddings provides richer user-item representations and thus improves recommendation accuracy.
@@
Despite the empirical success of these models, existing research has primarily concentrated on architectural innovations and fusion strategies (Liu et al., 2024; Zhou et al., 2023b; Fu et al., 2024b; Ge et al., 2024), while largely neglecting a more fundamental issue—the contribution of multimodal embeddings themselves to the observed performance gains. In particular, a large number of multimodal recommendation models have introduced complex mechanisms such as graph-based modeling of modalities and intricate fusion architectures (Zhuang et al., 2025). While these models are significantly more sophisticated than simple ID-based recommenders, if such complexity fails to yield tangible benefits, it raises a natural question: why use them at all? Although many works routinely integrate both text and image modalities, there is still a lack of comprehensive and controlled investigation into the actual utility of these embeddings. This presents a critical research gap. Motivated by this, we pose the central research question of this paper: Are multimodal embeddings truly beneficial for recommendation?
@@
Although extensive research has been conducted in the multimodal recommendation domain, evaluating the true effectiveness of multimodal embeddings remains challenging due to several factors (Beel et al., 2016): (1) most studies emphasize novel architectural designs or training strategies rather than directly examining the intrinsic quality of the embeddings; (2) evaluation protocols vary widely—including different datasets, metrics, and data-splitting strategies—making cross-study comparisons largely inconclusive; (3) reported results often reflect a combination of influences such as model architecture, optimization techniques, and training heuristics, which obscures the isolated contribution of the embeddings themselves.
@@ 2. Related Work
Reproducibility Study of Recommender Systems In the field of recommender systems, differences in datasets, evaluation protocols, and experimental settings make reproducibility studies particularly important (Beel et al., 2016). Rather than proposing new methodologies, these studies re-examine widely adopted baselines to draw fundamental insights for future research. For example, Ferrari Dacrema et al. (2019) reveals that many state-of-the-art neural network-based recommendation methods perform even worse than simple nearest-neighbor techniques, Rendle et al. (2020) revisits matrix factorization in comparison with neural collaborative filtering, and Petrov and Macdonald (2022) conducts a reproducibility study on the BERT4Rec model (Sun et al., 2019). Ji et al. (2023) presents a critical study on the offline evaluation of recommender systems. In the recent modality-based recommendation domain, many works (Li et al., 2023; Yuan et al., 2023; Fu et al., 2024c; Ni et al., 2023; Zhang et al., 2024a) compare advanced sequential recommendation models with heavy end-to-end encoders against simple ID embedding models, demonstrating that ID-based models still achieve strong performance. Although these studies do not propose novel methodologies, they provide valuable insights that have had a significant impact on the research community. Each serves as a cornerstone of the field, inspiring our study on recent state-of-the-art multimodal recommendation models.
@@
Multimodal recommendation. Multimodal recommendation (Liu et al., 2024; Zhou et al., 2023b; Fu et al., 2024a) leverages auxiliary information from different modalities, primarily text, and images, to enrich item representations. Early work like VBPR (He and McAuley, 2016) extended the classic Bayesian Personalized Ranking (BPR (Rendle et al., 2009)) framework by incorporating pre-extracted visual features into the matrix factorization model, demonstrating that visual style can influence user preference. With the rise of Graph Neural Networks (GNNs), which excel at capturing high-order connectivity, graph-based models became the dominant paradigm. MMGCN (Wei et al., 2019) was a pioneering work that constructed separate user-item bipartite graphs for each modality, learning modality-specific representations through graph convolution. Following this, GRCN (Wei et al., 2020) proposed refining the user-item graph by using multimodal features to identify and down-weight potential noisy (false-positive) interactions. More recent models like LATTICE (Zhang et al., 2021) and FREEDOM (Zhou and Shen, 2023) introduced an explicit item-item semantic graph built from multimodal features to capture latent item relationships, which is then combined with the user-item interaction graph. LGMRec (Guo et al., 2024) further extended this by learning both local and global user interests, using a hypergraph to model global dependencies. Other approaches, such as BM3 (Zhou et al., 2023a), adopt a self-supervised learning paradigm. Its core innovation lies in bootstrapping latent representations by generating contrastive views through a simple dropout mechanism in the embedding space, thereby eliminating the need for explicit and computationally expensive graph augmentation or construction.
@@
Concurrent work. Concurrent to ours, Pomo et al. (Pomo et al., 2025) examine whether multimodal recommendation gains stem from semantics or capacity, introducing LVLM(Large Vision–Language Model)-derived, multimodal-by-design embeddings via structured VQA prompting with EOS-based extraction. Meanwhile, Zhou et al. (Zhou et al., 2025) propose an evaluation framework and benchmark multimodal recommendation models across tasks, stages, and integration strategies to assess modality contributions. By contrast, our work diagnoses reliance via Modality Knockout: we replace the target modality with constants/noise during both training and inference while preserving the original model architecture, isolating data-signal effects and enabling consistent, model-agnostic comparison.
@@ 3. Preliminaries
where $(\cdot,\cdot)$ indicates a combination of the visual and textual embeddings, rather than a dimensional concatenation. The combined embedding $\tilde{\mathbf{e}}_{m_{i}}$ is then fed into a recommendation model, which may integrate it with user and item ID embeddings through various strategies such as simple fusion (e.g., VBPR (He and McAuley, 2016)), incorporation into a user–item interaction graph (e.g., MMGCN (Wei et al., 2019)), or by constructing an item–item similarity graph (e.g., FREEDOM (Zhou and Shen, 2023)). In this work, we dissect the contribution of each modality by analyzing their impact from two perspectives:
  1. (1)
@@
Modality Knockout. To rigorously answer our research question, we adopt a simple yet effective diagnostic methodology we named “Modality Knockout”. The central idea is to systematically replace the information from specific modalities and measure the resulting impact on the model’s final performance. Our method is inspired by (He et al., 2022), which examines scenarios where only a small subset of image patches is utilized to assess their representational capability, while the remaining patches are replaced with random noise.
@@
4. Experimental Setup
4.1. Datasets
We conduct our experiments on three widely-used Amazon (Ni et al., 2019) datasets, which are standard benchmarks in multimodal recommendation research, including Baby, Clothing, and Sports. Each dataset provides user–item interactions, item metadata including textual descriptions, and item images. We follow the standard data preprocessing steps used in the MMRec library [^1] https://github.com/enoche/MMRec/ to ensure fair comparison, including 5-core filtering and splitting data into training, validation, and test sets.
Table 1. Statistics of the Datasets.
@@ 4.1. Datasets
Item image features were obtained by using the pre-extracted vectors extracted by ResNet-50 (He et al., 2016) provided in the benchmark; these serve as the baseline visual embeddings. Item text features typically come from product titles or descriptions, encoded via pre-trained all-MiniLM-L6-v2 model (Reimers and Gurevych, 2020).
4.2. Models Implemented via the MMRec Toolbox
@@
    Simple fusion approaches: Traditional approaches integrate multimodal data in a straightforward manner—without graph-based fusion—by combining ID embeddings with other modalities via simple concatenation or direct contrastive learning.
  2. (2)
@@
    User–item graph integration: A representative work in this category is the popular MMGCN (Wei et al., 2019), which constructs separate user–item graphs for each modality to leverage multimodal signals.
  3. (3)
@@
    Item–item graph construction: Another paradigm exploits multimodal embeddings to build item-item graphs, capturing relationships between items through their multimodal content.
  4. (4)
@@
    Hybrid graph approaches: Another line of research combines the advantages of the above strategies, incorporating multimodal information into both user-item and item-item graph construction, thereby enhancing representation learning from both perspectives.
To provide a comprehensive evaluation, we conduct a large-scale empirical study across these four commonly adopted paradigms. Table 2 presents the selected models and their corresponding group assignments. A more detailed introduction of the adopted models is provided below:
Table 2. Grouping Multimodal Recommendation Models. We classify models by their use of multimodal embeddings for either User–Item (U-I) Graph integration (Int.) or Item–Item (I-I) construction (Cons.).
Model Group U-I Int. I-I Graph Cons. Models
Group 1 No No VBPR (He and McAuley, 2016), BM3 (Zhou et al., 2023a)
Group 2 Yes No GRCN (Wei et al., 2020), DualGNN (Wang et al., 2021), SLMRec (Tao et al., 2022), MMGCN (Wei et al., 2019), LGMRec (Guo et al., 2024)
Group 3 No Yes LATTICE (Zhang et al., 2021), FREEDOM (Zhou and Shen, 2023), MGCN (Yu et al., 2023), DA-MRS (Xv et al., 2024), SMORE (Ong and Khong, 2025)
Group 4 Yes Yes DRAGON (Zhou et al., 2023c), PGL (Yu et al., 2025)
4.2.1. Group 1: Simple Fusion Approaches
(1) VBPR (He and McAuley, 2016) is a foundational model in this category which extends the classic Bayesian Personalized Ranking (BPR) framework. It enhances item representations by incorporating pre-trained visual features, projecting them into the same latent space as the item ID embeddings, and adding a visual-specific preference term to the final ranking score. (2) BM3 (Zhou et al., 2023a) introduced a self-supervised learning approach. Its core innovation is bootstrapping latent representations by creating contrastive views through a simple dropout mechanism in the embedding space, thereby avoiding explicit and computationally expensive graph augmentation or construction.
4.2.2. Group 2: User-Item Graph Integration
(1) MMGCN (Wei et al., 2019) is a pioneering GNN-based model which constructs separate U-I graphs for each modality (e.g., visual, textual). It performs graph convolutions independently on each to learn modality-specific user preferences before fusing the resulting embeddings in the final layer. (2) GRCN (Wei et al., 2020) uses multimodal content as a signal to identify and prune potentially noisy edges in the U-I graph, thus cleaning the collaborative signal before propagation. (3) DualGNN (Wang et al., 2021) enhances the U-I graph by introducing a user-user co-occurrence graph, enabling the model to learn personalized modality fusion patterns by leveraging signals from similar users. (4) SLMRec (Tao et al., 2022) is a framework that applies self-supervised learning objectives to multimodal data to enhance recommendation performance, applying contrastive learning on augmented views of the U-I graph to learn more robust representations. (5) LGMRec (Guo et al., 2024) advances this paradigm by learning local user interests from the U-I graph while simultaneously modeling global interests through a hypergraph structure, which captures higher-order, attribute-based relationships to enrich user representations.
4.2.3. Group 3: Item-Item Graph Construction
(1) LATTICE (Zhang et al., 2021) is a foundational model in this group, which dynamically learns a latent I-I graph by fusing multiple modality-aware graphs constructed from item features. (2) FREEDOM (Zhou and Shen, 2023) is an evolution of LATTICE, improves efficiency by pre-calculating and ”freezing” the I-I graph structure before training begins. (3) MGCN (Yu et al., 2023) builds an I-I graph but first employs a behavior-guided purifier to denoise the raw multimodal features, ensuring that the constructed graph is based on high-quality, preference-aware signals. (4) SMORE (Ong and Khong, 2025) introduces a novel signal processing perspective, proposing to fuse and denoise multimodal features in the frequency domain before they are used in graph learning, which includes constructing I-I graphs to capture semantic correlations. (5) DA-MRS (Xv et al., 2024) is a model that utilizes disentangled representation learning and adversarial training to separate user intents and item attributes across modalities.
4.2.4. Group 4: Hybrid Graph Approaches
(1) DRAGON (Zhou et al., 2023c) explicitly constructs and learns from a U-I heterogeneous graph, a user-user graph, and an I-I graph derived from content similarity. It then fuses information from all three structures to generate comprehensive dual representations for users and items. (2) PGL (Yu et al., 2025) is also categorized as a hybrid approach. While its primary innovation is to combat over-smoothing by extracting a ”principal subgraph” from the U-I graph, its advanced graph learning mechanism implicitly leverages higher-order connectivity patterns that reflect latent item-item relationships, thus benefiting from both collaborative and semantic structural information.
4.3. Experimental Protocol
@@
To investigate the role of multimodal embeddings, we employ Modality Knockout as introduced in section 3 and define four experimental conditions. For each knockout setting, we consider both constant-value and random-noise knockouts. [^3] For constant-value knockouts, we primarily use ones as the default, since using zeros as the default can be incompatible with certain model implementations and may cause training to fail. Detailed results are provided in the Appendix Table 11. Specifically, we include these four settings in the experiments:
  1. (1)
@@
All models are trained to converge on the training set, and performance is reported on the test set using standard top-K recommendation metrics (He et al., 2017) including Recall and NDCG, and Precision.
Figure 1. Comparing 14 Multimodal Recommendation Models under Modality Knockouts on Baby, Clothing, and Sports Datasets.
Figure 2. Comparing Noise-Based and Constant Embedding Replacement with Baseline with Dual Knockout.
5. Main Experiment Results
@@ 5.1. Multimodality as a Whole (RQ1)
(Answer to RQ1) This demonstrates that the joint use of visual and textual features as a whole provides crucial information that substantially improves recommendation performance in most multimodal recommendation models with sophisticated fusion. Interestingly, common baselines with simpler fusion strategies, such as VBPR and BM3, exhibit surprisingly ”stable” performance, suggesting that the incorporated modality information is not beneficial in these models.
@@
Figure 3. Visualization of Text and Image Embeddings.
5.2. Individual Modalities (RQ2)
Having established the overall utility of multimodality, we now investigate the individual roles of the visual and textual modalities. We compare the results from the Visual Knockout and Textual Knockout experiments. The results reveal a striking asymmetry in how models depend on these two modalities. As demonstrated in subsection 5.1, the two knockout strategies exhibit similar trends. Due to space constraints, we present only the results for the random noise–based knockout, with the full experimental results provided in the Appendix A.
5.2.1. The Dominance of Text:
The Textual Knockout experiment, where only text embeddings were replaced, triggered a severe performance drop across almost all models, comparable to the drop seen in the Dual Knockout experiment. For example, on the Baby dataset, FREEDOM’s NDCG@10 dropped by 28.0%, DRAGON’s fell by 22.2% and LGMRec’s fell by 18.4%. This clearly indicates that models are heavily reliant on the textual modality. The rich semantic information contained in item titles and descriptions appears to be the primary driver of performance gains in most modern architectures.
5.2.2. The Ineffectiveness of Vision:
In stark contrast, the Visual Knockout experiment, where only image embeddings were replaced, had a minimal effect on the majority of models. In some cases, such as PGL on the Clothing dataset, removing the visual information even led to a slight performance increase. This strongly suggests that for a large portion of state-of-the-art models, the visual modality is severely underutilized, contributing little to the final recommendation. The models are not effectively “seeing” the images.

### v3

@@ Abstract
To isolate the effect of individual modalities—text or visual—we employ a modality knockout strategy by setting the corresponding embeddings to either constant values or random noise. To ensure the scale and comprehensiveness of our study, we evaluate 14 widely used state-of-the-art multimodal recommendation models. Our findings reveal that: (1) multimodal embeddings generally enhance recommendation performance—particularly when integrated through more sophisticated graph-based fusion models. Surprisingly, commonly adopted baseline models with simple fusion schemes, such as VBPR and BM3, show only limited gains. (2) The text modality alone achieves performance comparable to the full multimodal setting in most cases, whereas the image modality alone does not. These results offer foundational insights and practical guidance for the multimodal recommendation community. Our code and datasets are available under https://github.com/GAIR-Lab/MKF4MMRec.
1 Introduction
@@
Multimodal recommendation has emerged as a powerful paradigm in modern recommender systems, enabling the integration of diverse modalities such as text, image, video, and audio. By leveraging pretrained encoders such as Sentence-BERT, Vision Transformers (ViT), and ResNet, recent state-of-the-art models—including BM3 [39], FREEDOM [38], and SMORE [18]—have demonstrated impressive performance across various recommendation tasks. These models are built on the intuitive yet widely accepted assumption that incorporating multimodal embeddings provides richer user-item representations and thus improves recommendation accuracy.
@@
Despite the empirical success of these models, existing research has primarily concentrated on architectural innovations and fusion strategies [15, 36, 4, 7], while largely neglecting a more fundamental issue—the contribution of multimodal embeddings themselves to the observed performance gains. In particular, a large number of multimodal recommendation models have introduced complex mechanisms such as graph-based modeling of modalities and intricate fusion architectures [40]. While these models are significantly more sophisticated than simple ID-based recommenders, if such complexity fails to yield tangible benefits, it raises a natural question: why use them at all? Although many works routinely integrate both text and image modalities, there is still a lack of comprehensive and controlled investigation into the actual utility of these embeddings. This presents a critical research gap. Motivated by this, we pose the central research question of this paper: Are multimodal embeddings truly beneficial for recommendation?
@@
Although extensive research has been conducted in the multimodal recommendation domain, evaluating the true effectiveness of multimodal embeddings remains challenging due to several factors [1]: (1) most studies emphasize novel architectural designs or training strategies rather than directly examining the intrinsic quality of the embeddings; (2) evaluation protocols vary widely—including different datasets, metrics, and data-splitting strategies—making cross-study comparisons largely inconclusive; (3) reported results often reflect a combination of influences such as model architecture, optimization techniques, and training heuristics, which obscures the isolated contribution of the embeddings themselves.
@@ 2 Related Work
Reproducibility Study of Recommender Systems In the field of recommender systems, differences in datasets, evaluation protocols, and experimental settings make reproducibility studies particularly important [1]. Rather than proposing new methodologies, these studies re-examine widely adopted baselines to draw fundamental insights for future research. For example, [2] reveals that many state-of-the-art neural network-based recommendation methods perform even worse than simple nearest-neighbor techniques, [23] revisits matrix factorization in comparison with neural collaborative filtering, and [19] conducts a reproducibility study on the BERT4Rec model [24]. [13] presents a critical study on the offline evaluation of recommender systems. In the recent modality-based recommendation domain, many works [14, 32, 6, 17, 33] compare advanced sequential recommendation models with heavy end-to-end encoders against simple ID embedding models, demonstrating that ID-based models still achieve strong performance. Although these studies do not propose novel methodologies, they provide valuable insights that have had a significant impact on the research community. Each serves as a cornerstone of the field, inspiring our study on recent state-of-the-art multimodal recommendation models.
@@
Multimodal recommendation. Multimodal recommendation [15, 36, 3] leverages auxiliary information from different modalities, primarily text, and images, to enrich item representations. Early work like VBPR [11] extended the classic Bayesian Personalized Ranking (BPR [22]) framework by incorporating pre-extracted visual features into the matrix factorization model, demonstrating that visual style can influence user preference. With the rise of Graph Neural Networks (GNNs), which excel at capturing high-order connectivity, graph-based models became the dominant paradigm. MMGCN [28] was a pioneering work that constructed separate user-item bipartite graphs for each modality, learning modality-specific representations through graph convolution. Following this, GRCN [27] proposed refining the user-item graph by using multimodal features to identify and down-weight potential noisy (false-positive) interactions. More recent models like LATTICE [34] and FREEDOM [38] introduced an explicit item-item semantic graph built from multimodal features to capture latent item relationships, which is then combined with the user-item interaction graph. LGMRec [8] further extended this by learning both local and global user interests, using a hypergraph to model global dependencies. Other approaches, such as BM3 [39], adopt a self-supervised learning paradigm. Its core innovation lies in bootstrapping latent representations by generating contrastive views through a simple dropout mechanism in the embedding space, thereby eliminating the need for explicit and computationally expensive graph augmentation or construction. TARec [40] and FDRec [41] adopt Wasserstein-based knowledge distillation for multimodal recommendation. IISAN [3, 4] and CROSSAN [5] adopt a Decoupled PEFT structure for multimodal adaptation, matching full fine-tuning performance with reduced GPU memory usage.
@@
Concurrent work. Concurrent to ours, Pomo et al. [20] examine whether multimodal recommendation gains stem from semantics or capacity, introducing LVLM(Large Vision–Language Model)-derived, multimodal-by-design embeddings via structured VQA prompting with EOS-based extraction. Meanwhile, Zhou et al. [35] propose an evaluation framework and benchmark multimodal recommendation models across tasks, stages, and integration strategies to assess modality contributions. By contrast, our work diagnoses reliance via Modality Knockout: we replace the target modality with constants/noise during both training and inference while preserving the original model architecture, isolating data-signal effects and enabling consistent, model-agnostic comparison.
@@ 3 Preliminaries
where $(\cdot,\cdot)$ indicates a combination of the visual and textual embeddings, rather than a dimensional concatenation. The combined embedding $\tilde{\mathbf{e}}_{m_{i}}$ is then fed into a recommendation model, which may integrate it with user and item ID embeddings through various strategies such as simple fusion (e.g., VBPR [11]), incorporation into a user–item interaction graph (e.g., MMGCN [28]), or by constructing an item–item similarity graph (e.g., FREEDOM [38]). In this work, we dissect the contribution of each modality by analyzing their impact from two perspectives:
  1. 1.
@@
Modality Knockout. To rigorously answer our research question, we adopt a simple yet effective diagnostic methodology we named “Modality Knockout”. The central idea is to systematically replace the information from specific modalities and measure the resulting impact on the model’s final performance. Our method is inspired by [9], which examines scenarios where only a small subset of image patches is utilized to assess their representational capability, while the remaining patches are replaced with random noise.
@@
4 Experimental Setup
4.1 Datasets
Table 1: Statistics of the Datasets.
@@ 4.1 Datasets
We conduct our experiments on three widely-used Amazon [16] datasets, which are standard benchmarks in multimodal recommendation research, including Baby, Clothing, and Sports. Each dataset provides user–item interactions, item metadata including textual descriptions, and item images. We follow the standard data preprocessing steps used in the MMRec library [^1] https://github.com/enoche/MMRec/ to ensure fair comparison, including 5-core filtering and splitting data into training, validation, and test sets.
Item image features were obtained by using the pre-extracted vectors extracted by ResNet-50 [10] provided in the benchmark; these serve as the baseline visual embeddings. Item text features typically come from product titles or descriptions, encoded via pre-trained all-MiniLM-L6-v2 model [21].
4.2 Models Implemented via the MMRec Toolbox
@@
    Simple fusion approaches (Group 1): Traditional approaches integrate multimodal data in a straightforward manner—without graph-based fusion—by combining ID embeddings with other modalities via simple concatenation or direct contrastive learning. This group includes VBPR [11] and BM3 [39].
  2. 2.
@@
    User–item graph integration (Group 2): A representative work in this category is the popular MMGCN [28], which constructs separate user–item graphs for each modality to leverage multimodal signals. This group includes GRCN [27], DualGNN [26], SLMRec [25], MMGCN [28], and LGMRec [8].
  3. 3.
@@
    Item–item graph construction (Group 3): Another paradigm exploits multimodal embeddings to build item-item graphs, capturing relationships between items through their multimodal content. This group includes LATTICE [34], FREEDOM [38], MGCN [30], DA-MRS [29], and SMORE [18].
  4. 4.
@@
    Hybrid graph approaches (Group 4): Another line of research combines the advantages of the above strategies, incorporating multimodal information into both user-item and item-item graph construction, thereby enhancing representation learning from both perspectives. This group includes DRAGON [37] and PGL [31].
4.3 Experimental Protocol
@@
To investigate the role of multimodal embeddings, we employ Modality Knockout as introduced in section 3 and define four experimental conditions. For each knockout setting, we consider both constant-value and random-noise knockouts. [^3] For constant-value knockouts, we primarily use ones as the default, since using zeros as the default can be incompatible with certain model implementations and may cause training to fail. Specifically, we include these four settings in the experiments:
  1. 1.
@@
All models are trained to converge on the training set, and performance is reported on the test set using standard top-K recommendation metrics [12] including Recall and NDCG, and Precision.
Figure 1: Comparing 14 Multimodal Recommendation Models under Modality Knockouts on Baby, Clothing, and Sports Datasets.
Figure 2: Comparing Noise-Based and Constant Embedding Replacement with Baseline with Dual Knockout.
5 Main Experiment Results
@@ 5.1 Multimodality as a Whole (RQ1)
(Answer to RQ1) This demonstrates that the joint use of visual and textual features as a whole provides crucial information that substantially improves recommendation performance in most multimodal recommendation models with sophisticated fusion. Interestingly, common baselines with simpler fusion strategies, such as VBPR and BM3, exhibit surprisingly "stable" performance, suggesting that the incorporated modality information is not beneficial in these models.
@@
Figure 3: Visualization of Text and Image Embeddings.
5.2 Individual Modalities (RQ2)
Having established the overall utility of multimodality, we now investigate the individual roles of the visual and textual modalities. We compare the results from the Visual Knockout and Textual Knockout experiments. The results reveal a striking asymmetry in how models depend on these two modalities. As demonstrated in subsection 5.1, the two knockout strategies exhibit similar trends. Due to space constraints, we present only the results for the random noise–based knockout.
The Dominance of Text: The Textual Knockout experiment, where only text embeddings were replaced, triggered a severe performance drop across almost all models, comparable to the drop seen in the Dual Knockout experiment. For example, on the Baby dataset, FREEDOM’s NDCG@10 dropped by 28.0%, DRAGON’s fell by 22.2% and LGMRec’s fell by 18.4%. This clearly indicates that models are heavily reliant on the textual modality. The rich semantic information contained in item titles and descriptions appears to be the primary driver of performance gains in most modern architectures.
The Ineffectiveness of Vision: In stark contrast, the Visual Knockout experiment, where only image embeddings were replaced, had a minimal effect on the majority of models. In some cases, such as PGL on the Clothing dataset, removing the visual information even led to a slight performance increase. This strongly suggests that for a large portion of state-of-the-art models, the visual modality is severely underutilized, contributing little to the final recommendation. The models are not effectively “seeing” the images.